# Four-Plane Factorized Video Autoencoders

### v1

@@ 1 Introduction
One effective strategy to make generative modeling computationally feasible is through latent modeling, which enables generative models to operate on lower-dimensional representations [3, 38, 5, 34, 11, 20, 35, 18, 1]. Latent models use autoencoders to compress high-resolution visual data into a more compact latent space, where the generative process can proceed with a reduced computational burden – the generative model’s memory footprint and training and inference times are all directly improved by working in the smaller latent space.
Figure 1: Factorized latent representation. Traditional volumetric latents in diffusion models yield a sequence length of $t\times h\times w$ (top row), which scales linearly with the input size and demands high computational resources. Our proposed factorized representation reduces sequence length to $t\times(h+w)+2\times h\times w$ (bottom row), achieving a more compact latent space that scales sublinearly with input size, enabling faster, more efficient video generation; while maintaining high quality.
However, in typical autoencoders, the resulting latent size still scales linearly with the original input size, so the compression offers only a limited benefit when deployed in high-dimensional domains, such as handling videos with high spatial and temporal resolutions [20, 35] (Figure 1).
@@
  • The four-plane factorized autoencoder offers a significant reduction in latent resolution, as its size scales sublinearly with the total input size (see Figure 1). At the same time, signals decoded from the factorized latents are of comparable fidelity to those reconstructed from larger volumetric latents.
  • Our factorized autoencoder presents an effective latent representation for generative tasks – for example a typical transformer-based diffusion model used on volumetric features [20] generates samples twice as fast when trained on our factorized space, while maintaining comparable generation quality.
  • The latent structure is easily adaptable to a variety of image-conditioned generation tasks. We demonstrate this property on two common video applications, two-frame interpolation and frame extrapolation.
@@ 2.1 Diffusion models for video synthesis
Of the more recent diffusion models developed for video generation, many operate on a volumetric spatiotemporal latent space. VDM [24] employs a 3D U-Net autoencoder architecture [10, 39] to learn this space. The factorized space-time attention mechanism enables joint training on both images and videos. This architecture illustrates that handling spatial and temporal information simultaneously leads to more coherent video sequences.
To address scalability for high resolution video generation, Imagen Video [23] extends VDM by introducing a cascade of models that essentially alternate temporal and spatial superresolution. Lumiere [2] introduced the STUNet architecture, which generates entire videos directly with improved temporal coherence. VideoLDM [4] constructs a video model starting with pretrained image models and inserting temporal layers before fine-tuning on high-quality videos.
@@ 2.2 Video tokenizers
Many of the latent video diffusion models highlighted above rely on some form of video tokenization to compress high dimensional videos into a compact latent space. The pioneering vector quantization approaches for image tokenization, for example VQ-VAE [46], VQ-VAE2 [36], and VQGAN [15], can be applied to videos in a frame-by-frame manner. MAGVIT [51] introduced a 3D-VQ autoencoder to quantize video data into spatio-temporal tokens, making it a powerful tool for a range of video generation tasks such a frame prediction, video inpainting, and frame interpolation. MAGVIT-v2 [52] introduced significant advancements, including a lookup-free quantization method, which allows for an expanded vocabulary without compromising performance. Additionally, MAGVIT-v2 enables joint image and video modeling through a causal 3D CNN architecture. The MAGVIT-v2 tokenizer was used successfully in W.A.L.T [20] for photorealistic image and video generation. The early 3D encoder layers in our factorized architecture are based on MAGVIT-v2.
@@ 2.3 Video frame interpolation
Video frame interpolation [30, 14] has distinct interpretations depending on the temporal distance between frames. In the case where there is significant motion between frames, the challenging task can be addressed by generative models. The application of our factorized latent representation to frame interpolation can be categorized as a two-frame conditioned diffusion model. The first such effort to use a latent diffusion model was LDMVFI [12]. In contrast, ViDiM [28] models in pixel space and generates the entire video at once improving temporal consistency. To improve interpolation quality, ViDiM employs a cascaded diffusion approach: it first generates a low-resolution video, followed by an upsampling diffusion model that refines the output to higher resolutions
@@ 2.4 Tri-plane factorization
Tri-plane representations, which factorize volumetric data into three orthogonal 2D planes, have been widely used as compact representations of 3D neural fields [42, 7, 16]. Coupled with diffusion model architectures suited to planar representations, they have been used for a variety of applications, such as textured 3D model generation [50], 3D neural field generation [42], and semantic scene generation [31]. The same concept was applied to videos in PVDM [54], where encoding videos to tri-plane latent features enables the use of a 2D U-Net architecture [13] for training diffusion models, bringing significant improvements in efficiency. In contrast, our work proposes a distinctive four-plane factorization approach, which broadens the model’s scope, enabling tasks like frame extrapolation and video interpolation.
Figure 2: Model overview. The tokenizer first maps the input video into a volumetric latent representation through $3D$ convolutional architecture, which is then factorized into four planes. Temporal planes are created by mean pooling along the height and width dimensions, capturing time-varying features. Spatial planes are obtained by splitting along the time axis and independently averaging along this dimension, focusing on spatial consistency (highlighted in green). During decoding, the four planes are mapped back into a volume: for each spatial-temporal location, features from the corresponding four planes (highlighted in blue) are concatenated to reconstruct the full volumetric features. These combined features are passed through a decoder to produce the final output video.
@@ 3 Factorized video latent representations
In latent video diffusion models, a key component is the autoencoder, which compresses the input video data into a compact latent representation. To achieve this, prior works [20, 21] employ a 3D convolutional architecture that encodes the 3D video volume $\mathbf{X}\in\mathbb{R}^{T\times H\times W\times C}$ into a feature volume $\mathbf{Z}\in\mathbb{R}^{t\times h\times w\times c}$, where $t=\frac{T}{f_{t}}$, $h=\frac{H}{f_{s}}$, $w=\frac{W}{f_{s}}$. Here $f_{t}$ and $f_{s}$ are the temporal and spatial downsampling factors. The channel dimension $c$ is typically expanded (e.g., $c=8$ is a common choice of channel dimension that balances autoencoder reconstruction and diffusion performance).
@@ 3.1 Four plane factorization
Our factorization approach is intended to directly and simply reduce the cubic footprint of volumetric spatiotemporal latent spaces. At a high level, our approach decomposes the 3D feature volume into four distinct planes, each capturing complementary aspects of the video’s spatial and temporal features. This leads to a more efficient representation that accelerates training while preserving the necessary information for high-quality reconstruction. Our introduction of a four-plane factorization is motivated by the versatility it brings, making it adaptable to a range of tasks such as class-conditional generation, frame extrapolation, and video interpolation.
@@
Given an input video $\mathbf{X}\in\mathbb{R}^{T\times H\times W\times 3}$, our encoder network first converts it into a feature volume $\mathbf{Z}\in\mathbb{R}^{t\times h\times w\times c}$ using a causal 3D convolution architecture similar to the one introduced in MAGVIT-v2 [53]. The feature volume is then factorized into four planes along three directions: two spatial planes, $\mathbf{P}_{xy}^{1},\mathbf{P}_{xy}^{2}\in\mathbb{R}^{h\times w\times c}$, aligned along the $xy$-dimension, and two spatiotemporal planes, $\mathbf{P}_{xt}\in\mathbb{R}^{t\times h\times c}$ and $\mathbf{P}_{yt}\in\mathbb{R}^{t\times w\times c}$, aligned along the $xt$ and $yt$ dimensions, respectively.
To obtain the latent feature planes, we apply a factorization operation along certain dimensions. Specifically, the two spatio-temporal planes $\mathbf{P}_{xt}$ and $\mathbf{P}_{yt}$ are obtained by applying a factorization operation along the height and width dimensions, respectively:
@@
where $\Lambda_{h}$ and $\Lambda_{w}$ performs pooling or dimensionality reduction across spatial dimensions $h$ and $w$ resulting in the planar representation.
Similarly, the spatial planes contain the temporally aggregated features across frames, capturing the spatial structure and background information in the video. To obtain the spatial planes $\mathbf{P}_{xy}^{1}$ and $\mathbf{P}_{xy}^{2}$, we adapt our factorization approach based on the application. For class-conditional generation and frame prediction, we split the latent feature volume $\mathbf{Z}\in\mathbb{R}^{t\times h\times w\times c}$ into two segments along the temporal dimension, then aggregate each segment over $T$, yielding:
@@
[…]
(UCF dataset) (Kinetics-600 dataset)
Video Diffusion [24] - 16.2 1.1B 256
RIN [27] - 10.7 411M 1000
TATS [17] 332 - 321M 1024
Phenaki [48] - 36.4 227M 48
MAGVIT [51] 76 9.9 306M 12
MAGVIT-v2 [53] 58 4.3 307M 24
W.A.L.T [20] 46 TML]B3E5FC3.3 313M 50
W.A.L.T* [20] TML]E3F6FF39 TML]E3F6FF5.7 313M 50
Ours TML]B3E5FC38 8.6 214M 50
Table 1: Class-conditional generation on UCF and frame prediction on Kinetics-600. WALT* represents our re-training and re-evaluation of the WALT baseline. Our method achieves competitive performance with WALT on the UCF task and performs slightly lower on the K600 frame prediction task, showcasing efficient performance across both datasets.
For the interpolation task, where the first and last frames are available, we leverage this information by using a pre-trained image tokenizer to directly obtain the latent representations of these frames. Specifically, we set:
@@
where $E$ denotes the pretrained image tokenizer and $\mathbf{X}_{0}$ and $\mathbf{X}_{T}$ are the boundary frames. This approach effectively incorporates key frame information into the spatial planes, enhancing the model’s interpolation accuracy.
@@ 3.1.2 Recomposition
where $\mathrm{Combine}$ represents the chosen operation, e.g., $\mathrm{Combine}(\cdot)=\sum$ for summation, or $\mathrm{Combine}(\cdot)=\mathrm{concat}(\cdot)$ for concatenation, along the channel dimension.
This reconstructed feature volume $\mathbf{V}$ is then fed into a decoder with a structure similar to MAGVIT-v2, which progressively upsamples the features and applies 3D convolutions to generate the final video $\hat{\mathbf{X}}\in\mathbb{R}^{T\times H\times W\times 3}$.
@@ 3.2 Generative modeling with factorized latents
Latent diffusion models (LDMs) gradually transform the latent representation of data into noise in a forward diffusion process, then reverse this transformation to generate new samples. Given an initial latent sample $\mathbf{x}_{0}$, the forward process generates a sequence of increasingly noisy latents $\{\mathbf{z}_{t}\}_{t=1}^{T}$ as follows:
$\mathbf{z}_{t}=\sqrt{\alpha_{t}}\mathbf{z}_{t-1}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\mathbf{x}_{0}\approx D(\mathbf{z}_{0}),$ (5)
where $\alpha_{t}$ controls the noise schedule and $D$ decodes the final latent back to data space. The reverse process, defined by $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})=\mathcal{N}(\mathbf{z}_{t-1};\mu_{\theta}(\mathbf{z}_{t},t),\sigma_{t}^{2}\mathbf{I})$, seeks to denoise the latent variables and reconstruct the original data distribution. This denoising is learned by minimizing a variational lower bound, often simplified into practical objectives [22]. Here, we adopt the v-parameterization, following recent diffusion improvements [41].
To train a LDM on our factorized representation we use a transformer architecture similar to W.A.L.T. [20]. We create a 1D sequence by flattening the four planes—$\mathbf{P}_{xt}$, $\mathbf{P}_{yt}$, $\mathbf{P}^{1}_{xy}$, and $\mathbf{P}^{2}_{xy}$—and concatenating them along the sequence length dimension. This results in a sequence length of $h\times t+w\times t+2\times h\times w$ (as shown in Figure 1).
@@ 4 Experiments
In order to evaluate the performance of the proposed four-plane factorized autoencoder, we combine it with a diffusion model (Section 3.2) and apply it towards three different experimental benchmarks: class-conditional video generation, future frames prediction, and two-frame interpolation. The observed performance on these diverse synthesis experiments demonstrates the versatility of the proposed four-plane latent space.
Figure 3: Class-conditional generation results on the UCF dataset. The model generates 17 frames at a resolution of 256 x 256. Here, we display 9 frames, selected by subsampling every other frame (stride of 2), from W.A.L.T and our model, to illustrate the continuity and quality of generated video sequences.

### v2

# Factorized Video Autoencoders for Efficient Generative Modelling

@@ 1 Introduction
One effective strategy to make generative modeling computationally feasible is through latent modeling [3, 41, 5, 37, 10, 19, 38, 17, 1]. By compressing high-resolution visual data into a compact latent space, latent models significantly reduce the computational burden for generative models. However, in typical latent model autoencoders, the resulting embedding size still scales linearly with the original input size, so this compression offers only a limited benefit when deployed in very high dimensional domains, such as videos [19, 38] (see Figure 1).
Figure 1: Factorized latent representation. Traditional volumetric latents in diffusion models yield a sequence length of $t\times h\times w$ (top row), which scales linearly with the input size and demands high computational resources. Our proposed factorized representation reduces sequence length to $t\times(h+w)+2\times h\times w$ (bottom row), achieving a more compact latent space that scales sublinearly with input size, enabling faster, more efficient video generation without sacrificing quality.
@@
  • Compact yet expressive factorization: four-plane factorization significantly compresses volumetric latent embeddings, scaling sublinearly with the total input size (see Fig. 1). Despite this compression, it retains high-fidelity reconstructions.
  • Faster generation without sacrificing quality: we demonstrate the factorized latent space is suitable for high quality generative modeling. A conventional transformer-based diffusion model trained to generate factorized latents is twice as fast compared to producing volumetric features [19].
  • Versatility for image-conditioned tasks: our experiments show how the four-plane structure seamlessly accommodates a variety of applications such as two-frame interpolation and future frame prediction.
@@ 2.1 Diffusion models for video synthesis
Of the more recent diffusion models developed for video generation, many operate on a volumetric spatiotemporal latent space. VDM [23] employs a 3D U-Net autoencoder architecture [9, 42] to learn this space. To address scalability for high resolution video generation, Imagen Video [22] extends VDM by introducing a cascade of models that essentially alternate temporal and spatial superresolution. Lumiere [2] introduced the STUNet architecture, which generates entire videos directly with improved temporal coherence. VideoLDM [4] constructs a video model starting with pretrained image models and inserting temporal layers before fine-tuning on high-quality videos.
@@ 2.2 Video tokenizers
Many of the latent video diffusion models highlighted above rely on some form of video tokenization [19, 62, 55, 15, 52] to compress high dimensional videos into a compact latent space. The pioneering vector quantization approaches for image tokenization, for example VQ-VAE [49], VQ-VAE2 [39], and VQGAN [14], can be applied to videos in a frame-by-frame manner. MAGVIT [56] introduced a 3D-VQ autoencoder to quantize video data into spatio-temporal tokens, making it a powerful tool for a range of video generation tasks such a frame prediction, video inpainting, and frame interpolation. MAGVIT-v2 [57] introduced significant advancements, including a lookup-free quantization method, which allows for an expanded vocabulary without compromising performance. Additionally, MAGVIT-v2 enables joint image and video modeling through a causal 3D CNN architecture. The MAGVIT-v2 tokenizer was used successfully in W.A.L.T [19] for photorealistic image and video generation. The early 3D encoder layers in our factorized architecture are based on MAGVIT-v2.
@@ 2.3 Video frame interpolation
Video frame interpolation [29, 13] has distinct interpretations depending on the temporal distance between frames. In case with significant motion between frames, the challenging task can be addressed by generative models. The application of our factorized latent representation to frame interpolation can be categorized as a two-frame conditioned diffusion model. The first such effort to use a latent diffusion model was LDMVFI [11]. In contrast, VIDIM [27] models in pixel space and generates the entire video at once improving temporal consistency. To improve quality, VIDIM employs a cascaded diffusion approach: it first generates a low-resolution video, followed by an upsampling diffusion model that refines the output to higher resolutions
@@ 2.4 Tri-plane factorization
Tri-plane representations, which factorize volumetric data into three orthogonal 2D planes, have been widely used as compact representations of 3D neural fields [45, 7, 16]. Coupled with diffusion model architectures suited to planar representations, they have been used for a variety of applications, such as textured 3D model generation [54], 3D neural field generation [45], and semantic scene generation [32]. The same concept was applied to videos in PVDM [59], where encoding videos to tri-plane latent features enables the use of a 2D U-Net architecture [12] for training diffusion models, bringing significant improvements in efficiency. However, tri-plane approaches are still quite far behind their volumetric counterparts in generation quality, and the tri-plane representation cannot support different use cases. Both of these points are addressed by our distinctive four-plane factorization.
Figure 2: Model overview. The autoencoder first maps the input video into a volumetric latent representation through $3D$ convolutional architecture, which is then factorized into four planes. Temporal planes are created by mean pooling along the height and width dimensions, capturing time-varying features. Spatial planes are obtained by splitting along the time axis and independently averaging along this dimension, focusing on spatial consistency (highlighted in green). During decoding, the four planes are mapped back into a volume: for each spatial-temporal location, features from the corresponding four planes (highlighted in blue) are concatenated to reconstruct the full volumetric features. These combined features are passed through a decoder to produce the final output video.
@@ 3 Factorized video latent representations
In latent video diffusion models, a key component is the autoencoder, which compresses the input video data into a compact latent representation. To achieve this, prior works [19, 20] employ a 3D convolutional architecture that encodes the 3D video volume $\mathbf{X}\in\mathbb{R}^{T\times H\times W\times C}$ into a feature volume $\mathbf{Z}\in\mathbb{R}^{t\times h\times w\times c}$, where $t=\frac{T}{f_{t}}$, $h=\frac{H}{f_{s}}$, $w=\frac{W}{f_{s}}$. Here $f_{t}$ and $f_{s}$ are the temporal and spatial downsampling factors. The channel dimension $c$ is typically expanded (e.g., $c=8$ is a common choice that balances autoencoder reconstruction and diffusion performance).
@@ 3.1 Four plane factorization
Our approach provides a direct and effective solution to reducing the cubic complexity of volumetric spatiotemporal latent spaces. By decomposing the 3D feature volume into four complementary planes, our method captures rich spatial and temporal structures while significantly improving efficiency. This streamlined representation not only accelerates training and inference but also reduces memory overhead without compromising reconstruction quality. Beyond efficiency gains, the four-plane factorization introduces a versatile framework adaptable to a wide range of video generation tasks, including class-conditional generation, frame extrapolation, and video interpolation.
@@
Given an input video $\mathbf{X}\in\mathbb{R}^{T\times H\times W\times 3}$, our encoder network first converts it into a feature volume $\mathbf{Z}\in\mathbb{R}^{t\times h\times w\times c}$ using a causal 3D convolution architecture similar to the one introduced in MAGVIT-v2 [58]. The feature volume is then factorized into four planes along three directions: two spatial planes, $\mathbf{P}_{xy}^{1},\mathbf{P}_{xy}^{2}\in\mathbb{R}^{h\times w\times c}$, aligned along the $xy$-dimension, and two spatiotemporal planes, $\mathbf{P}_{xt}\in\mathbb{R}^{t\times h\times c}$ and $\mathbf{P}_{yt}\in\mathbb{R}^{t\times w\times c}$, aligned along the $xt$ and $yt$ dimensions, respectively. The two spatio-temporal planes $\mathbf{P}_{xt}$ and $\mathbf{P}_{yt}$ are obtained by collapsing the height and width dimensions, respectively:
@@
where $\Lambda_{h}$ and $\Lambda_{w}$ performs pooling or dimensionality reduction operation along the required dimensions.
The spatial planes $\mathbf{P}_{xy}^{1}$ and $\mathbf{P}_{xy}^{2}$ contain the temporally aggregated features across frames, capturing the spatial structure and background information in the video. We adapt our factorization approach based on the application. For class-conditional generation and frame prediction, we split the latent feature volume $\mathbf{Z}\in\mathbb{R}^{t\times h\times w\times c}$ into two segments along the temporal dimension, then aggregate each segment over $T$, yielding:
@@
For the interpolation task only the first and last frames are available at inference and therefore the spatial plane cannot contain information from other timesteps. To address this, we obtain the spatial planes by encoding only the first and last frames using our encoder. Specifically, we set:
@@
where $E$ denotes our video encoder and $\mathbf{X}_{0}$ and $\mathbf{X}_{T}$ are the boundary frames. Since our model uses a 3D CNN with causal temporal padding, it can naturally encode images without requiring additional modifications. This approach effectively incorporates key frame information into the spatial planes, enhancing the model’s interpolation accuracy.
@@ 3.1.2 Recomposition
where $\mathrm{Combine}$ represents the recomposition operation, e.g., summation or concatenation, along the channel dimension.
The reconstructed feature volume $\mathbf{V}$ is then fed into a decoder with a structure similar to MAGVIT-v2, which progressively upsamples the features and applies 3D convolutions to generate the final video $\hat{\mathbf{X}}\in\mathbb{R}^{T\times H\times W\times 3}$.
@@ 3.2 Generative modeling with factorized latents
Latent diffusion models (LDMs) gradually transform the latent representation of data into noise in a forward diffusion process, then reverse this transformation to generate new samples. Given an initial latent sample $\mathbf{x}_{0}$, the forward process generates a sequence of increasingly noisy latents $\{\mathbf{z}_{t}\}_{t=1}^{T}$ as $\mathbf{z}_{t}=\sqrt{\alpha_{t}}\mathbf{z}_{t-1}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\mathbf{x}_{0}\approx D(\mathbf{z}_{0})$ where $\alpha_{t}$ controls the noise schedule and $D$ decodes the final latent back to data space. The reverse process, defined by $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})=\mathcal{N}(\mathbf{z}_{t-1};\mu_{\theta}(\mathbf{z}_{t},t),\sigma_{t}^{2}\mathbf{I})$, seeks to denoise the latent variables and reconstruct the original data distribution. This denoising is learned by minimizing a variational lower bound, often simplified into practical objectives [21]. Here, we adopt the v-parameterization, following recent diffusion improvements [44].
To train a LDM on our factorized representation we use a transformer architecture. We create a 1D sequence by flattening the four planes—$\mathbf{P}_{xt}$, $\mathbf{P}_{yt}$, $\mathbf{P}^{1}_{xy}$, and $\mathbf{P}^{2}_{xy}$—and concatenating them along the sequence length dimension. This results in a sequence length of $h\times t+w\times t+2\times h\times w$ (as shown in Figure 1).
@@ 4 Experiments
The focus of our experiments is to show that the four-plane factorized model can generally and seamlessly replace volumetric latent spaces when modeling videos. To understand the model’s representation capability, we will quantify the compression versus reconstruction tradeoff (Sec. 4.1). To demonstrate its widespread applicability, we deploy the factorized latent space in varied generative tasks such as class-conditional video generation (Sec. 4.2), future frames prediction (Sec. 4.3), and two-frame interpolation (Sec. 4.4). Notably, the factorized autoencoder design is used without modification across experiments, as only the training data changes to ensure fair comparisons for each task.
4.1 Autoencoder reconstruction
One primary consideration of the proposed factorization is if the considerable latent compression comes at a cost in representation quality. For this, we compare the reconstructions from a baseline volumetric autoencoder against our factorized approach. The factorized encoder and decoder share the same architecture as the volumetric baseline, differing only in the factorization and recomposition steps. Specifically, to obtain the planes from the latent feature volume, we use average pooling as the factorization operations $\Lambda_{h}$ , $\Lambda_{w}$ and $\Lambda_{t}$. To recompose the planes into feature volume, we use the concatenation operation as $\mathrm{Combine}$. We discuss further architecture and training details in the apendix.
We chose the W.A.L.T. [19] autoencoder as our volumetric baseline model as it has shown state of the art performance on multiple benchmark tasks such as class-conditional video generation and frame extrapolation. Furthermore we were able to reproduce the model in terms of similar datasets and performance, allowing us to evaluate it in new settings as well.
We trained the volumetric and our four-plane factorized model on the Kinetics-600 (K600) [6] dataset, consisting of nearly 400,000 video clips, covering around 600 action classes and exhibiting a wide range of human activities.  Tab. 1 shows the reconstruction quality for $128\times 128$ resolution videos is nearly identical for both the volumetric and factorized autoencoders, while the latent size (sequence length) is nearly halved for the factorized. We also train a $256\times 256$ model by attaching an extra layer to the encoder and decoder architecture (see appendix for details), leaving the latent size unchanged and observe comparable performance for the volumetric and fourplane autoencoder.
Whereas W.A.L.T. adopts a continuous autoencoder framework, some recent works for latent video generation have opted for a variational autoencoder (VAE [31]) design [52, 55, 62]. Our factorized proposal is not predicated on this decision, so we would expect similar conclusions for our approach in the VAE setting. To validate this, we conduct two additional experiments. First, we modify W.A.L.T. autoencoder to incorporate a VAE decoder and construct a corresponding factorized-VAE variant. Second, we adopt WF-VAE [34] from OpenSoraPlan [35], training the VAE using our factorized representation and comparing its performance against the baseline. For experiments with the W.A.L.T. autoencoder, we use a latent dimensionality of 8, while for WF-VAE, we set the latent dimensionality to 4.  Table 1 confirms the relative performance holds across both VAE and AE settings. Since W.A.L.T. (AE) is the established baseline, our subsequent evaluations are all performed in the AE setting.
[…]
128 Volumetric 27.64 0.85 0.049 1280
4Plane 27.11 0.82 0.051 672
Volumetric-VAE 27.32 0.84 0.053 1280
4Plane-VAE 27.03 0.82 0.055 672
256 Volumetric 26.27 0.79 0.089 1280
4Plane 25.67 0.77 0.104 672
WF-VAE 27.86 0.83 0.064 1280
4Plane-WF-VAE 26.98 0.81 0.073 672
Table 1: Video reconstruction. We show reconstruction metrics for our four-plane model and the volumetric baseline (W.A.L.T. [19]) on the Kinetics-600 test set, alongside the corresponding sequence length induced by each latent representation. We evaluate tokenizers trained using both autoencoder and VAE frameworks.